OpenStudy (anonymous):

Please help! will fan and medal! How would i put this into the proper format? 16x^2 + 25y^2 + 32x - 150y = 159

OpenStudy (anonymous):

it needs to be the proper form of an elipse.

OpenStudy (anonymous):

\[\left( 16 x^2+32x \right)+\left( 25y^2-150 y \right)=159\] 1.complete the squares 2. make right hand side =1 by dividing the number on R.H.S

OpenStudy (anonymous):

so \[(4x+\sqrt{32x}+5y-\sqrt{150y})/159\]

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

use first step first ,then solve by second step.

OpenStudy (anonymous):

x - 4 = \frac{1}{4}(y + 1)^2 goes to \[\frac{ 16x^2+32x }{ 159 }+\frac{ 25y^2-150y }{ 159}=1\]?

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

\[y=a(x-h)^2+k\]

OpenStudy (anonymous):

\[16\left( x^2+2x+1-1 \right)+25\left( y^2-6y+9-9 \right)=159\] \[16\left( x+1 \right)^2-16+25\left( y-3 \right)^2-225=159\] \[16(x+1)^2+25\left( y-3 \right)^2=159+16+225=400\] divide by 400 \[\frac{ \left( x+1 \right)^2 }{ 25 }+\frac{ \left( y-3 \right)^2 }{ 16 }=1\]

OpenStudy (anonymous):

thank you sorry I accidentally started looking at another problem and confused myself.

OpenStudy (anonymous):

yw